AWildDove:

The graph below shows the height through which an elevator travels, y, in x seconds: A graph is shown with x axis title as Time in seconds. The title on the y axis is Height of Elevator Above Ground in feet. The values on the x axis are from 0 to 5 in increments of 1 for each grid line. The values on the y axis are from 0 to 125 in increments of 25 for each grid line. A line is shown connecting points on ordered pair 1, 25 and 2, 50 and 3, 75 and 4, 100. The title of the graph is Rate of Ascent. What is the rate of change for the relationship represented in the graph?

AWildDove:

Forgot the anwsers 50 25 fraction 1 over 25 fraction 1 over 50

Imagine:

So your function box should look like this y+1| 1 | 25 | x+25 | 2 | 50 | | 3 | 75 | | 4 | 10 |

AWildDove:

Yes

Imagine:

And get rid of anything that says 50, because your value goes up 25.

AWildDove:

Right I have i think its 25 but im not too sure

Imagine:

And It is not 25/1 so, It has to be y=1/25=x

AWildDove:

Okay so it would be choice C correct?

Imagine:

Correct.

AWildDove:

Thank you!

Imagine:

You're welcome!

Imagine:

wait. Justus can you make sure I'm correct?

AWildDove:

@justus ?

justus:

Sure!

justus:

The rate of change is just like speed. 50-25/2-1=25

justus:

So it is not the third option.

AWildDove:

So would it be the 2nd option?

justus:

Yes I believe it is the 2nd option.

AWildDove:

Okay thank you!

justus:

No problem.